The sum of two integers is twenty. Five times the smaller integer is two more than twice the larger integer. Find the two integers.
Answer by ptaylor(2198)

Let x=smaller integer
And 20-x=the larger integer
Now we are told that:
5*x-2=2(20-x) get rid of parens
5x-2=40-2x add 2 and also 2x to each side
5x-2+2+2x=40+2-2x+2x collect like terms
7x=42 divide each side by 7
x=6-------------------------------smaller integer
20-x=20-6=14 larger integer


CK
5*6-2=2*14
28=28
This could also be solved using two unknowns
x=smaller
y=larger
x+y=20-------------------eq1
5x-2=2y--------------------eq2
Hope this helps---ptaylor